Deen Enterprises currently sells its products for $1600 per unit. Management is contemplating a 20% increase in the selling price for the next year. Variable costs are currently 30% of sales revenue and are not expected to change next year. Fixed expenses are $280,800 per year.

What is the breakeven point in units at the anticipated selling price per unit next year?

480 units

195 units

1755 units

117 units
Answer #1

Current variable cost = 1600*30% = 480

New sales price = 1600 + 20% = 1920

Breakeven point

= Fixed cost /Contribution margin per unit

= 280,800/(1920 - 480)

= 195 units

Option B is the answer